OpenStudy (anonymous):

Part 1: Solve the following system of equations. Please show your work to receive full credit. x - y = 10 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

Hellppp!!!

OpenStudy (anonymous):

x=4 and y=-6

OpenStudy (anonymous):

can you show ur work plz?

OpenStudy (anonymous):

If you add both the equations you can easily remove the y , x-y=10 2x+y=2(Add them) ------- 3x=12 x=4, put value of x in any one of the equation let it be used in the first equation 4-y=10 y=-6